# http://www.michaeltimothybennett.com/

### v4

@@ Abstract
The concept of intelligent software is flawed. The behaviour of software depends upon the hardware that interprets it. This undermines claims regarding the behaviour of theorised, software superintelligence. Here we characterise this problem as “computational dualism”, where instead of mental and physical substance, we have software and hardware. We argue that to make objective claims regarding performance we must avoid computational dualism. We propose using an alternative based upon pancomputationalism, wherein every aspect of the environment is a relation between irreducible states. We formalise systems as behaviour (inputs and outputs), and cognition as embodied, embedded, extended and enactive. The result is cognition formalised as a part of the environment, rather than as a disembodied policy interacting with the environment though an interpreter. This allows us to make objective claims regarding intelligence, which we argue is the ability to “generalise”, identify causes and adapt. We then propose objective upper bounds for intelligent behaviour.
@@ 2 Formalising Enactivism
Embodiment: We use these dyadic relations to formalise enactivism [thompson2007], holding that everything we call the mind is just part of the environment, and “thinking” is just changes in environmental state. This blurs the line between agent and environment, making the distinction unclear. As Heidegger maintained, Being is bound by context [heidegger2020]. There is no need to define an agent that has no environment, and so there seems to be little point in preserving the distinction. What we need is not a Cartesian model of disembodied intelligence looking in upon an environment but an embodiment, embedded in a particular part of the environment, through which goal directed behaviour can be enacted. We need to describe the circuitry with which cognition is embodied and enacted, much like a formal language only with truth conditions determined by environmental states. Just as every computational system we can build is finite, and living systems are ergodic [friston2013], we assume all of this takes place within a system which has a finite number of configurations, and so amounts to a finite number of declarative programs. We call this an abstraction layer. Intuitively, an abstraction layer is like a window. It looks onto that part of the environment in which cognition takes place. For example, we might enumerate every possible declarative program which pertains only to one specific computer, and use that computer as our abstraction layer. However, as we can take any set of declarative programs and define an abstraction layer, it is “pancomputational” in the sense that it captures every system in computational terms.
@@ 3 Limits of Intelligence
We have equated intelligence with sample efficient generalisation. According to [bennett2023b, prop. 1,2] the weakest correct policies have the highest probability of generalising. Given an uninstantiated task $\lambda_{\rho}$, utility to measures the weakness of the weakest correct policies in different vocabularies. By choosing $\mathfrak{v}$ which maximises utility for $\lambda_{\rho}(\mathfrak{v})$, we instantiate $\lambda_{\rho}$ in a vocabulary that maximises the weakness of correct policies for $\lambda_{\rho}$. Then, using weakness proxy, we can select a policy that has the highest possible probability of generalising, and thus maximise sample efficiency. ∎
Put another way, utility is maximised for $\lambda_{\rho}(\mathfrak{v})$ when $\mathfrak{v}=P$. It follows that when utility is maximised there must exist $\pi\in\Pi_{\lambda_{\rho}(\mathfrak{v})}$ which is a weakest policy in $\Pi_{\lambda_{\rho}(P)}$. This is why maximising utility maximises the weakness of correct policies.
Concluding remarks: Here we proposed computational dualism, and presented an argument in support of an alternative which allows for objective claims regarding behaviour. We proposed an objective upper bound for intelligent behaviour. In general this upper bound is unattainable, because we can only build systems with finite vocabularies. Moreover, too large a vocabulary $\mathfrak{v}$ could make $\mathfrak{v}$-tasks intractable. However in practical terms higher utility does mean weaker and more generalise-able policies, which suggests one should optimise for higher utility whilst trying to minimise the size of the vocabulary. We anticipate these results may further our understanding of AI safety and general intelligence.

### v7

@@ Abstract
The concept of intelligent software is flawed. The behaviour of software is determined by the hardware that “interprets” it. This undermines claims regarding the behaviour of theorised, software superintelligence. Here we characterise this problem as “computational dualism”, where instead of mental and physical substance, we have software and hardware. We argue that to make objective claims regarding performance we must avoid computational dualism. We propose a pancomputational alternative wherein every aspect of the environment is a relation between irreducible states. We formalise systems as behaviour (inputs and outputs), and cognition as embodied, embedded, extended and enactive. The result is cognition formalised as a part of the environment, rather than as a disembodied policy interacting with the environment through an interpreter. This allows us to make objective claims regarding intelligence, which we argue is the ability to “generalise”, identify causes and adapt. We then establish objective upper bounds for intelligent behaviour. This suggests AGI will be safer, but more limited, than theorised.
@@ 2 Formalising Enactivism
Embodiment: We use these dyadic relations to formalise enactivism [thompson2007], holding that everything we call the mind is just part of the environment, and “thinking” is just changes in environmental state. This blurs the line between agent and environment, making the distinction unclear. As Heidegger maintained, Being is bound by context [heidegger2020]. There is no need to define an agent that has no environment, and so there seems to be little point in preserving the distinction. What we need is not a Cartesian model of disembodied intelligence looking in upon an environment but an embodiment, embedded in a particular part of the environment, through which goal directed behaviour can be enacted. We need to describe the circuitry with which cognition is embodied and enacted, much like a formal language only with truth conditions determined by whatever laws govern the environment. Just as every computational system we can build is finite, and living systems are ergodic [friston2013], we assume all of this takes place within a system which has a finite number of configurations, and so amounts to a finite number of declarative programs. We call this an abstraction layer. Intuitively, an abstraction layer is like a window. It looks onto that part of the environment in which cognition takes place. For example, we might enumerate every possible declarative program which pertains only to one specific computer, and use that computer as our abstraction layer. However, as we can take any set of declarative programs and define an abstraction layer, it is “pancomputational” in the sense that it captures every system in computational terms.
@@ 3 Limits of Intelligence
We have equated intelligence with sample efficient generalisation. According to [bennett2023b, prop. 1,2] the weakest correct policies have the highest probability of generalising. Given an uninstantiated task $\lambda_{\rho}$, utility measures the weakness of the weakest correct policies. We can use this to compare vocabularies. By choosing a vocabulary $\mathfrak{v}$ which maximises utility for $\lambda_{\rho}(\mathfrak{v})$, we instantiate $\lambda_{\rho}$ in a vocabulary that maximises the weakness of correct policies for $\lambda_{\rho}$. Then, using weakness proxy, we can select a policy that has the highest possible probability of generalising, and thus maximise sample efficiency. ∎
Put another way, utility is maximised for $\lambda_{\rho}(\mathfrak{v})$ when $\mathfrak{v}=P$. When utility is maximised there must exist $\pi\in\Pi_{\lambda_{\rho}(\mathfrak{v})}$ which is a weakest policy in $\Pi_{\lambda_{\rho}(P)}$.
Concluding remarks: Here we argued software minds are a flawed concept, symptomatic of “computational dualism”. We argued for an alternative which allows for objective claims regarding behaviour. We used this alternative to propose upper bounds on intelligence, and we anticipate these results may further our understanding of AI safety and AGI. In practical terms this upper bound is unattainable, because we can only build systems with finite vocabularies. Moreover, too large a vocabulary $\mathfrak{v}$ could make $\mathfrak{v}$-tasks intractable. Physical embodiment severely constrains what is possible. Rigorous research has been undertaken regarding the risks of AGI [cohen2022], but it is based upon computational dualism. Our results suggest AGI will be safer, but more limited, than has been theorised. Our results may also be of use in the pursuit of AGI. Higher utility does mean weaker and more generalise-able policies, which suggests one should optimise for higher utility whilst trying to minimise $\lvert\mathfrak{v}\rvert$.